gababb09:

An equation is shown below: 4(2x - 5) + 15 = 11 Write the steps you will use to solve the equation and explain each step.

TutoriaIs:

First, do you know the rules of PEMDAS?

TutoriaIs:

P-parenthesis, E-Exponents, M-Multiplication, D-Division, A-Addition, S-Subtraction

snowflake0531:

The distributive property says \(a(b-c) = a(b)-a(c)\) So the first step would be to distribute the parenthesis \(4(2x - 5) + 15 = 11\\4(2x)-4(5)+15=11\\8x-20+15=11\) From there you want to add the like terms \(8x-5=11\) Then you want to isolate the variable \(8x-5+5=11+5\\8x=16\) Then make it 1x \(\frac{8x}{8}=\frac{16}{8}\\x=2\)

gababb09:

Thanks for helping

snowflake0531:

yw~